OpenStudy (anonymous):

What does asymptotic behavior mean? I am asked to find that but I'm not sure what they are looking for. How do you write one?

OpenStudy (anonymous):

For example, what would be the asymptotic behavior of \[\frac{ x }{ x^2-x-2 }\]?

OpenStudy (phi):

when x = -1 the bottom becomes -1*-1 - (-1) -2 = 1 + 1 -2 = 2-2=0 dividing by 0 is not allowed... however if let x get very close to -1, the bottom gets very close to 0 (it is a very tiny number) when you divide by a tiny number you get a big number. Example: 1/ 0.001 = 1000 (use a calculator if you don't see this) as the bottom gets closer to 0 (gets tinier) you divide by a tinier number and get a bigger number. the numbers get bigger as x gets closer to -1, and zoom up to infinity you never get to x =-1 but you asymptote to x=-1

OpenStudy (anonymous):

hmmm I get what you are saying, but is there a specific way to write asymptotic behavior?

OpenStudy (anonymous):

yes \[\lim_{x\to \infty }\frac{ x }{ x^2-x-2 }=0\]